OpenStudy (anonymous):

How to solve m/n=p-6 for n

OpenStudy (unklerhaukus):

multiply both sides of equation by n, then divide both sides of equation by (p-6)

OpenStudy (diyadiya):

\[\frac{m}{n}=p-6\]by multiplying both sides by n we get \[m=n(p-6)\]dividing both sides by (p-6) we get \[n=\frac{m}{(p-6)}\]

OpenStudy (anonymous):

wouldnt you get n/p-6 =m/p-6?

hero (hero):

You only had one (p-6) to begin with. How do you end up with two of them?

OpenStudy (diyadiya):